# 00footnotetext: Shared first author; Shared last author

### v1

@@ Appendix I Model-based quality filters
Results are contained in Table˜18. The main conclusions we can arrive at from this table are as follows: i) Suffix Array deduplication seems to help more than MinHash deduplication, thereby giving some signal to the source of the gains procured by a MinHash+SuffixArray pipeline; ii) BFF provides comparable performance to a full Exact+MinHash+SuffixArray pipeline, giving strong evidence that the multiresolution BFF could be an easily scalable alternative to the relatively more expensive MinHash+SuffixArray pipeline of prior works. Interestingly, it appears that a SuffixArray pipeline seems to outperform MinHash alone, though this falls within the range of variance for the Core score due to the nondeterminism in subsampling the dataset and training a model.
@@
Results are contained in Table˜19. The first point to note is that BFF with a min_ngram_size at 13 and 20 yields Core scores and MMLU scores that are comparable to the scores attained by a MinHash+SuffixArray deduplicated pool at the same scale. The second point to note regards the BFF min_ngram_size and sharding: interestingly a lower min_ngram_size yields higher Core scores, but lower MMLU scores. We also see that fewer shards decreases the token yield, but has variable effect on the Core score. We examine the hyperparameters for BFF more fully in the next subsection.
@@
Encouraged by these results, next we examine the top candidates for a scalable deduplication pipeline at the 7B-2x scale. Again we start with a pool obtained from one tenth of Common Crawl and generate several deduplicated pools. The questions of interest are the same as above and we summarize the results in Table˜20. The key takeaways from this round of ablations is that at the 7B-2x scale, BFF with a min_ngram_size of 13 and 10 shards attains nearly identical performance to a MinHash+SuffixArray pipeline, whereas BFF with a min_ngram_size of 20 and 32 shards starts to lag behind, and that a min_ngram_size of 5 yields competitive Core scores, but falters in MMLU evaluations. While these experiments also vary the sharding choice, we view sharding primarily as a choice made to trade-off scalability with token yield. Larger shards are more expensive and less parallelizable and can decrease the token yield. For this round of ablations, the primary interest is to gain signal about how BFF compares to MinHash and Suffix Arrays at scale, and which are the correct hyperparameters for BFF. On this latter point, we chose to move forward with a min_ngram_size of 13 for generating DCLM-baseline.
Table 20: Deduplication Ablations (7B-2x scale). From the same pools as in Table˜19, we trained and evaluated models at the 7B-2x scale. Notice that a min_ngram_size of 5 yields competitive Core results but drastically reduces MMLU scores.
@@
Here we start with the 75B token data pool as in Section˜K.2.1 and focus on a paragraph-only level BFF. In other words, we run BFF as described above, except omit the full-document removal step. We use the default hyperparameters for ngram sizes as in Groeneveld [65], of 5 and 13 for min_ngram_size and max_ngram_size and a threshold of 0.8. We specifically look at the effect of changing the false positive rate and compute the removal rate (in bytes) of the output. From Table˜21, we can see that a false positive rate of 0.1 suffices for a reasonably small pool such as this one. For larger pools, to be safe, we always set the false positive rate to 0.01.
@@
From Table˜19 and Table˜20, we saw that altering the ngram_size hyperparameters can affect both token yield and evaluation metrics. In particular, we seek to examine how surviving documents are altered by deduplication. As a proxy for this, we focus on the document lengths and removal rates. Results for this paragraph and the two following paragraphs are collated in Table˜22. One key observation is that as the min_ngram_size parameter is reduced, the mean and median document lengths become shorter. This indicates that too-low a min_ngram_size parameter can dramatically affect the language statistics of the dataset and should be avoided. This tracks with intuitive sense where many documents include linebreak separated lists where each list element is short and possibly repeated: e.g., many webpages include recipes that might call for "1 stick of butter", which would get removed with a min_ngram_size of 5 but would injuriously damage the source document.
@@
Finally we simulate how shards affect the statistics of the deduplicated datasets. As above, the key statistics we focus on here are the removal rate and the average and medium document lengths. This is mostly to get a sense for how these features change as the dataset scales, with the prevailing thought that dramatically altering document statistics might adversely effect downstream evaluations. For larger pools, we can always shard them as heavily as desired, so we treat sharding as a hyperparameter that controls removal rate and document statistics. Results are collated in Table˜23 and Figure˜7. The key takeaways here are that removal rates increase monotonically with dataset size as expected, but do so in a concave fashion. This provides some signal for how heavily to shard an input pool if a desired token yield is specified. The next point of interest is to consider the document lengths as the dataset scales. These decrease monotonically as the pool increases in size.
@@
In building DCLM-baseline, at the point of deduplication, the dataset is approximately 70TB in size. Since Table˜20 shows that BFF had the best performance at a 10-way shard with a roughly 7TB input size, we adhere to a 100-way sharding for DCLM-baseline, where each shard is roughly 700GB in size.
@@
Finally, to get a sense for the duplicates remaining in a dataset after a full processing pipeline has been applied, we run a global (i.e., one shard) MinHash on several open datasets. These results are collated in table table˜24. We evaluate our DCLM-baseline, the official RefinedWeb dataset from HuggingFace, our emulation of the RefinedWeb pipeline, and Dolma V1. MinHash is performed using 14 buckets and a bucket size of 9, corresponding to the green curve in fig.˜6.
@@
RefinedWeb (official) 968B MinHash+SA 1* 0%
@@
In Section˜4.5, we showed that mixing our dataset with the usual sources did not improve its general performance, and hypothesized that this is due to the more stringent filtering performed for our Common Crawl portion. One could argue that improved filtering in the other sources could lead to similar improvements in performance. As such, we perform an experiment where we apply the same fastText classifier for filtering the other sources, as we do for our DCLM-baseline.
We take several source from RedPajama [156], and filter them with the fastText classifier applied in our DCLM-baseline, while keeping only the highest scored ones. We then add the resulting data to our pretraining dataset, and train models at the 1B-1x scale. The results of this can be seen in Table˜25. We see that despite the more uniform handling of mixing across various sources, the additional sources still decrease performance.
@@
We then evaluate various quality filters from Section˜4.4 on this data to search for correlation between dataset quality (as measured by Core accuracy) and filter agreement with human labels. Figure˜8 (a) depicts the Core scores of models trained on datasets filtered with the respective quality filter against ROC-AUC [^5] ROC-AUC measures a classifier’s ability to distinguish between classes by summarizing the trade-off between true positive and false positive rates across all thresholds, making it a robust and common metric for model performance. See https://en.wikipedia.org/wiki/Receiver_operating_characteristic for further details. of our quality-filters on the Majority data. Notably, both the best and worst fastText-based filters score about the same on the Majority data ($\sim 73\%$ ROC-AUC), while the AskLLM filter that is highly correlated with human annotations ($\penalty 10000\ \sim 82\%$ ROC-AUC) performs much worse as a quality filter ($\sim 28.5\%$ Core compared to $31\%$ in several fastText classifiers).
We continue this study by inspecting correlations to specific downstream tasks and comparing them to the ROC-AUC on the Agreement data, where all three annotators agreed on the label. Figure˜8 depicts the scores on a few tasks against the ROC-AUC on the annotated data of the representative set of quality filters. While some positive correlation may be observed for StrategyQA [60], the opposite is true for other QA datasets such as SQuAD [130], and in both cases the $R^{2}<0.3$. In most other downstream tasks, the results are similar to Figure˜8 (a), where no correlation can be observed. This suggests that human intuition may not reliably identify the most useful documents for language model training purposes. We hypothesize that human curators may create datasets that lack sufficient diversity and leave further investigation of these hypotheses to future research.
@@
We sampled 499 random documents from our pool, after going through the rule-based quality filters and deduplication. Figure˜9 shows a histogram of the length of the documents in words. We asked 16 English speak AI graduate and professors to annotate each example as a good candidate to be included in an LM pretraining corpus (see instructions and some examples given to annotators below). [^6] All annotation efforts were done voluntarily and were not paid for. Out of the 499 samples, in 281 samples there was full agreement between all three annotators. We release both datasets in https://datacomp.ai/dclm.
@@
        {internallinenumbers*} Your task is to assess the quality of the documents (0-bad, 1-good), sampled randomly from RefinedWebV2, in terms of their usefulness for LLM training.              How do you judge the usefulness of a document?              While this is a subjective task, there are a few things to keep in mind:        {internallinenumbers*} 1. We encourage you to review the details of the evaluation tasks that the LLMs are expected to excel in [link to spreadsheet with the tasks was provided]. Broadly, these tasks include reading comprehension, language understanding, world knowledge, and commonsense reasoning.        {internallinenumbers*} 2. Check whether a particular document will be useful for answering any open-ended imperative tasks or questions. For example: writing a letter to a friend, writing a poem, replying to an email, cooking a dish, etc.        {internallinenumbers*} 3. Check whether the data contains harmful content that you would not want an LLM to generate. This can include personal information, vulgar language, etc.        4. See the examples below to build some intuition.
@@
        Host: Zander        Program Category: Music        Frequency: Weekly        Length: 2 Hours        Terms: Barter        Delivery Method: Internet        {internallinenumbers*} |“Zander’s knowledge of music and his straight-forward approach has struck a huge interest among our listeners. The Rockin’ 80’s is EXACTLY what we’ve been looking for!"" - Terry West, WQLA        {internallinenumbers*} The Rockin’ 80’s is the only 80’s show with a mix of the best rock from the decade of excess plus “oh wow” tracks that add spice to the weekly line up. The two hour version of the show features rarities from the 80’s “Lost and Found”, an 80’s “Two-Fer,” spotlighting two contrasting songs from one band played back to back and much more. Featured core musical artists include Guns N’ Roses, Motley Crue, Van Halen, Rush and AC/DC.        {internallinenumbers*} |3733 Park East Drive • Room 222 • Cleveland, Ohio 44122 P: 216-831-3761 • F: 216-514-4699 • Email us        |©2014 Envision Networks. All rights reserved.        Site design by Single Source Marketing
@@
        |Chinese Five-Spice Noodles with Shitake Mushrooms        |Recipes - Chinese Five Spice        Chinese Five-Spice Noodles with Shitake Mushrooms        {internallinenumbers*} Cook the noodles until tender according to the package directions. Drain and rinse under cold running water, and drain well again.        {internallinenumbers*} Bring 2 1/2 cups water to a boil in a small saucepan. Add the dried mushrooms, cover, and simmer for 3 minutes. Strain the liquid through a paper coffee filter into a bowl to remove any grit, then squeeze the mushrooms over the bowl. Roughly chop the mushrooms, then set aside with the liquid.        {internallinenumbers*} Heat the butter and oil in a wok or large saucepan over medium heat. Add the Chinese five-spice powder and shallots and saute, stirring frequently, until tender, about 2 minutes. Add the garlic and saute 2 minutes more. Increase the heat to medium-high and add the fresh and reconstituted-dried shitakes with 1/2 cup of the reserved mushroom water. Cook, stirring frequently, until the mushrooms are tender, about 3 minutes. Add the soy sauce and rice wine. Increase the heat to high and cook, stirring, for 1 minute. Add the remaining mushroom-soaking water.        {internallinenumbers*} Add the noodles to the wok and stir until heated through and coated with the sauce, about 1 minute. Garnish with the green onions and sesame seeds, if using, and serve at once.        Makes 4 servings
@@
        {internallinenumbers*} "Is it necessary to purchase a travel book or is it realistic that we can get similar information from other resources? Usually, most individuals have a major question on buying a travel book. So here are the pros and cons of purchasing one such book.        Advantages of a Travel Book        {internallinenumbers*} A travel book, which may be a paperback or e-book, comes in handy while traveling. Glancing through a travel book enables you to understand the custom and culture of a particular place in the world. So you can adapt yourself to that particular environment and stay there comfortably for longer periods.        {internallinenumbers*} - They Come In Handy — The travel guide comes in various forms such as, e-books, paperbacks and the file formats. You can have easy access to these books, which would assist you with all details compatible to the region you are traveling to.        {internallinenumbers*} - They Provide Enormous Information — Electronic or traditional travel guides provide you with answers to all types of questions such as how to learn some sayings that can be used in the place where you are traveling to? How to get data on where to reside, what to see and where to eat? How to get a clear knowledge about the history of a specific region or the atmosphere that it has?        {internallinenumbers*} - They Suit To Your Requirements — To access full information about a specific country or a region, both types of general and specific travel books are made available. The e-book may easily fit into your e-book reader whereas the paperback can fit into your backpack.        Disadvantages of Travel Book        {internallinenumbers*} - The Price — The e-book and paperback travel guides are very expensive compared to the information obtained from travel websites or from those who have moved or traveled to that region.        {internallinenumbers*} - Qualitative Images In Travel Books — Most travel books are in black and white. Only a few e-books consist of colored photos. Hence make a thorough revision before purchasing a travel guide or an e-book.        {internallinenumbers*} - Travel Books Make The Trip Less Natural — Traveling can be made more spontaneous by acquiring suggestions from locals than from travel books."
@@
In Section˜4.6, we examined the results of the contamination analysis with respect to MMLU. Here, we present some analysis with respect to the other validation sets. Instead of the MMLU-specific decontamination that we performed in Section˜4.6, here we follow a more general approach based on token overlaps.
@@
We examine the difference of performance of the same 7B-2x model trained on DCLM-baseline, between the full evaluation set and the evaluation samples that are marked as “not dirty” per the criterion in Touvron et al. [158] (less than 80 % of the tokens in the sample are marked as contaminated), and between the full evaluation set and samples marked as “clean” using the same criterion (less than 20 % of the tokens are marked). Results can be seen in Figure˜10, where we see that the difference in performance over the full dataset and the “not dirty” samples is minimal. In fact, for BoolQ and SQuAD, which are marked as highly contaminated, our model performs slightly better on the “not dirty” subset. Moreover, in Figure˜11 we see that the difference in performance between the full evaluation set and the “clean” subset is similarly small for most datasets. We note here that it’s difficult to identify a correct threshold for what counts as a contaminated sample (as 20 % token overlap might lead to many false positives, but at the same time 80 % might be too high to detect all contaminated samples).
@@
Instruction tuning has emerged as a critical step to allow users to interact with pretrained language models [164, 162, 108, 111, 179]. To investigate whether models trained on DCLM-baseline can have strong instruction-following capabilities, we instruction-tune the DCLM-baseline (7B) with the OpenHermes-2.5 (OH-2.5) dataset [153]. Specifically, we train our model to predict the response given the instruction from the instruction-tuning dataset. We train DCLM-baseline using the Adam [83] optimizer for $10$ epochs with a $10\%$ warmup ratio and a cosine learning rate schedule. We perform a hyperparameter search over the three learning rates $\{1\text{e-}7,5\text{e-}6,2\text{e-}5\}$ and report the best-performing numbers on the evaluation metrics i.e., AlpacaEval 2.0 length-controlled win-rate [48]. Post-training, we generate the responses for the instructions from AlpacaEval using a sampling temperature of $0.3$ and maximum response length of $500$. We benchmark our model with relevant baselines (e.g., LLaMA-2-Chat [158], Zephyr-7B [159]) taken directly from the AlpacaEval leaderboard as well as Mistral-7B [75] finetuned in the same manner as DCLM-baseline.
@@
We present the results in Table˜26. We find that DCLM-baseline finetuned with OH-2.5 outperforms various instruct models such as Zephyr-Beta-7B and Gemma-Instruct-7B. This indicates that we can elicit high-quality responses from the pretrained DCLM-baseline model with instruction-tuning. In addition, we observe that the DCLM-baseline slightly lags behind Mistral-7B-OH-2.5 meaning DCLM-baseline is competitive with other existing models of the same scale for finetuning. The small difference in performance might be attributed to the DCLM-baseline w/ OH-2.5 having longer generations on average than Mistral-7B w/ OH-2.5 or the lesser number of tokens seen during DCLM-baseline pretraining in comparison to Mistral-7B.
A follow-up question is whether DCLM-baseline can be finetuned to be even more competitive with models of similar scale. To further improve the instruction-following capabilities of DCLM-baseline, we curate a custom dataset, DCLM-IT, by combining some of the best instruction-tuning datasets including UltraFeedback [43], Tulu-v2 SFT [74], CodeFeedback [187], OH-2.5, Nectar [190], NoRobots [129], WildChat [185], WebInstruct [180], and StarCoder2-Self-OSS-Instruct [165]. There are roughly $4$ million instances and $8$ billion tokens in this dataset. Subsequently, we perform instruction-tuning and response generation of DCLM-baseline on this dataset with the training recipe mentioned above. We present the results in Table˜26. We find that DCLM-IT outperforms DCLM-baseline w/ OH-2.5 by $2.8$ percentage points. Our results highlight that there is room to enhance the instruction-following capabilities of DCLM-baseline with better datasets such as DCLM-IT. We further clarify that the current instruction-tuned models do not undergo any alignment procedures such as PPO [139], DPO [126] or others [50, 30, 15, 104]. We leave the development of aligned versions of DCLM-baseline for future research.
@@
The final run 2.5T shown in Figure˜1 and Table˜9 uses a two stage training procedure as followed in Groeneveld et al. [66], Hu et al. [72] and Team et al. [151]. For stage 1 we use the hyper-parameters from Table˜27.
@@
After 2T tokens, we cooldown on a re-weighted pre-training distribution. For the cooldown distribution we use a mix of 70% DCLM-baseline with a tighter fastText threshold (top 7% rather than top 10%) and 30% ProofPile. We keep all the hyperparameters the same as Table˜27, so we cooldown to the same final learning rate, just over a smaller number of tokens. Before the cool-down MMLU performance was approximately 52%, and the LR was approximately $1\times 10^{-3}$.
We performed 2 independent cooldowns, one for 270B tokens and another for 200B tokens, and created a “model soup" [169] with a weight of 0.8 on the 270B cooldown and a weight of 0.2 on the 200B cooldown. Thus we the total number of tokens seen by this model is 2.5T. We present results of each individual cooldown and the model soup in Table˜28. The model in Figure˜1 and Table˜9 uses the final model soup after long-context training for 100B tokens as described in Section˜P.2.

### v2

@@ Appendix I Model-based quality filters
Results are contained in Table 18. The main conclusions we can arrive at from this table are as follows: i) Suffix Array deduplication seems to help more than MinHash deduplication, thereby giving some signal to the source of the gains procured by a MinHash+SuffixArray pipeline; ii) BFF provides comparable performance to a full Exact+MinHash+SuffixArray pipeline, giving strong evidence that the multiresolution BFF could be an easily scalable alternative to the relatively more expensive MinHash+SuffixArray pipeline of prior works. Interestingly, it appears that a SuffixArray pipeline seems to outperform MinHash alone, though this falls within the range of variance for the Core score due to the nondeterminism in subsampling the dataset and training a model.
@@
Results are contained in Table 19. The first point to note is that BFF with a min_ngram_size at 13 and 20 yields Core scores and MMLU scores that are comparable to the scores attained by a MinHash+SuffixArray deduplicated pool at the same scale. The second point to note regards the BFF min_ngram_size and sharding: interestingly a lower min_ngram_size yields higher Core scores, but lower MMLU scores. We also see that fewer shards decreases the token yield, but has variable effect on the Core score. We examine the hyperparameters for BFF more fully in the next subsection.
@@
Encouraged by these results, next we examine the top candidates for a scalable deduplication pipeline at the 7B-2x scale. Again we start with a pool obtained from one tenth of Common Crawl and generate several deduplicated pools. The questions of interest are the same as above and we summarize the results in Table 20. The key takeaways from this round of ablations is that at the 7B-2x scale, BFF with a min_ngram_size of 13 and 10 shards attains nearly identical performance to a MinHash+SuffixArray pipeline, whereas BFF with a min_ngram_size of 20 and 32 shards starts to lag behind, and that a min_ngram_size of 5 yields competitive Core scores, but falters in MMLU evaluations. While these experiments also vary the sharding choice, we view sharding primarily as a choice made to trade-off scalability with token yield. Larger shards are more expensive and less parallelizable and can decrease the token yield. For this round of ablations, the primary interest is to gain signal about how BFF compares to MinHash and Suffix Arrays at scale, and which are the correct hyperparameters for BFF. On this latter point, we chose to move forward with a min_ngram_size of 13 for generating DCLM-baseline.
Table 20: Deduplication Ablations (7B-2x scale). From the same pools as in Table 19, we trained and evaluated models at the 7B-2x scale. Notice that a min_ngram_size of 5 yields competitive Core results but drastically reduces MMLU scores.
@@
Here we start with the 75B token data pool as in Section K.2.1 and focus on a paragraph-only level BFF. In other words, we run BFF as described above, except omit the full-document removal step. We use the default hyperparameters for ngram sizes as in Groeneveld [67], of 5 and 13 for min_ngram_size and max_ngram_size and a threshold of 0.8. We specifically look at the effect of changing the false positive rate and compute the removal rate (in bytes) of the output. From Table 21, we can see that a false positive rate of 0.1 suffices for a reasonably small pool such as this one. For larger pools, to be safe, we always set the false positive rate to 0.01.
@@
From Table 19 and Table 20, we saw that altering the ngram_size hyperparameters can affect both token yield and evaluation metrics. In particular, we seek to examine how surviving documents are altered by deduplication. As a proxy for this, we focus on the document lengths and removal rates. Results for this paragraph and the two following paragraphs are collated in Table 22. One key observation is that as the min_ngram_size parameter is reduced, the mean and median document lengths become shorter. This indicates that too-low a min_ngram_size parameter can dramatically affect the language statistics of the dataset and should be avoided. This tracks with intuitive sense where many documents include linebreak separated lists where each list element is short and possibly repeated: e.g., many webpages include recipes that might call for "1 stick of butter", which would get removed with a min_ngram_size of 5 but would injuriously damage the source document.
@@
Finally we simulate how shards affect the statistics of the deduplicated datasets. As above, the key statistics we focus on here are the removal rate and the average and medium document lengths. This is mostly to get a sense for how these features change as the dataset scales, with the prevailing thought that dramatically altering document statistics might adversely effect downstream evaluations. For larger pools, we can always shard them as heavily as desired, so we treat sharding as a hyperparameter that controls removal rate and document statistics. Results are collated in Table 23 and Figure 7. The key takeaways here are that removal rates increase monotonically with dataset size as expected, but do so in a concave fashion. This provides some signal for how heavily to shard an input pool if a desired token yield is specified. The next point of interest is to consider the document lengths as the dataset scales. These decrease monotonically as the pool increases in size.
@@
In building DCLM-baseline, at the point of deduplication, the dataset is approximately 70TB in size. Since Table 20 shows that BFF had the best performance at a 10-way shard with a roughly 7TB input size, we adhere to a 100-way sharding for DCLM-baseline, where each shard is roughly 700GB in size.
@@
Finally, to get a sense for the duplicates remaining in a dataset after a full processing pipeline has been applied, we run a global (i.e., one shard) MinHash on several open datasets. These results are collated in table table 24. We evaluate our DCLM-baseline, the official RefinedWeb dataset from HuggingFace, our emulation of the RefinedWeb pipeline, and Dolma V1. MinHash is performed using 14 buckets and a bucket size of 9, corresponding to the green curve in fig. 6.
@@
RefinedWeb (official) 968M MinHash+SA 1* 0%
@@
In Section 4.5, we showed that mixing our dataset with the usual sources did not improve its general performance, and hypothesized that this is due to the more stringent filtering performed for our Common Crawl portion. One could argue that improved filtering in the other sources could lead to similar improvements in performance. As such, we perform an experiment where we apply the same fastText classifier for filtering the other sources, as we do for our DCLM-baseline.
We take several source from RedPajama [160], and filter them with the fastText classifier applied in our DCLM-baseline, while keeping only the highest scored ones. We then add the resulting data to our pretraining dataset, and train models at the 1B-1x scale. The results of this can be seen in Table 25. We see that despite the more uniform handling of mixing across various sources, the additional sources still decrease performance.
@@
We then evaluate various quality filters from Section 4.4 on this data to search for correlation between dataset quality (as measured by Core accuracy) and filter agreement with human labels. Figure 8 (a) depicts the Core scores of models trained on datasets filtered with the respective quality filter against ROC-AUC [^5] ROC-AUC measures a classifier’s ability to distinguish between classes by summarizing the trade-off between true positive and false positive rates across all thresholds, making it a robust and common metric for model performance. See https://en.wikipedia.org/wiki/Receiver_operating_characteristic for further details. of our quality-filters on the Majority data. Notably, both the best and worst fastText-based filters score about the same on the Majority data ($\sim 73\%$ ROC-AUC), while the AskLLM filter that is highly correlated with human annotations ($\leavevmode\nobreak\ \sim 82\%$ ROC-AUC) performs much worse as a quality filter ($\sim 28.5\%$ Core compared to $>31\%$ in several fastText classifiers).
We continue this study by inspecting correlations to specific downstream tasks and comparing them to the ROC-AUC on the Agreement data, where all three annotators agreed on the label. Figure 8 depicts the scores on a few tasks against the ROC-AUC on the annotated data of the representative set of quality filters. While some positive correlation may be observed for StrategyQA [62], the opposite is true for other QA datasets such as SQuAD [133], and in both cases the $R^{2}<0.3$. In most other downstream tasks, the results are similar to Figure 8 (a), where no correlation can be observed. This suggests that human intuition may not reliably identify the most useful documents for language model training purposes. We hypothesize that human curators may create datasets that lack sufficient diversity and leave further investigation of these hypotheses to future research.
@@
We sampled 499 random documents from our pool, after going through the rule-based quality filters and deduplication. Figure 9 shows a histogram of the length of the documents in words. We asked 16 English speak AI graduate and professors to annotate each example as a good candidate to be included in an LM pretraining corpus (see instructions and some examples given to annotators below). [^6] All annotation efforts were done voluntarily and were not paid for. Out of the 499 samples, in 281 samples there was full agreement between all three annotators. We release both datasets in https://datacomp.ai/dclm.
@@
     Your task is to assess the quality of the documents (0-bad, 1-good), sampled randomly from RefinedWebV2, in terms of their usefulness for LLM training.    How do you judge the usefulness of a document?    While this is a subjective task, there are a few things to keep in mind:      1. We encourage you to review the details of the evaluation tasks that the LLMs are expected to excel in [link to spreadsheet with the tasks was provided]. Broadly, these tasks include reading comprehension, language understanding, world knowledge, and commonsense reasoning.      2. Check whether a particular document will be useful for answering any open-ended imperative tasks or questions. For example: writing a letter to a friend, writing a poem, replying to an email, cooking a dish, etc.      3. Check whether the data contains harmful content that you would not want an LLM to generate. This can include personal information, vulgar language, etc.      4. See the examples below to build some intuition.
@@
     Host: Zander  Program Category: Music  Frequency: Weekly  Length: 2 Hours  Terms: Barter  Delivery Method: Internet  |“Zander’s knowledge of music and his straight-forward approach has struck a huge interest among our listeners. The Rockin’ 80’s is EXACTLY what we’ve been looking for!"" - Terry West, WQLA  The Rockin’ 80’s is the only 80’s show with a mix of the best rock from the decade of excess plus “oh wow” tracks that add spice to the weekly line up. The two hour version of the show features rarities from the 80’s “Lost and Found”, an 80’s “Two-Fer,” spotlighting two contrasting songs from one band played back to back and much more. Featured core musical artists include Guns N’ Roses, Motley Crue, Van Halen, Rush and AC/DC.  |3733 Park East Drive • Room 222 • Cleveland, Ohio 44122 P: 216-831-3761 • F: 216-514-4699 • Email us  |©2014 Envision Networks. All rights reserved.  Site design by Single Source Marketing
@@
     |Chinese Five-Spice Noodles with Shitake Mushrooms  |Recipes - Chinese Five Spice  Chinese Five-Spice Noodles with Shitake Mushrooms  Cook the noodles until tender according to the package directions. Drain and rinse under cold running water, and drain well again.  Bring 2 1/2 cups water to a boil in a small saucepan. Add the dried mushrooms, cover, and simmer for 3 minutes. Strain the liquid through a paper coffee filter into a bowl to remove any grit, then squeeze the mushrooms over the bowl. Roughly chop the mushrooms, then set aside with the liquid.  Heat the butter and oil in a wok or large saucepan over medium heat. Add the Chinese five-spice powder and shallots and saute, stirring frequently, until tender, about 2 minutes. Add the garlic and saute 2 minutes more. Increase the heat to medium-high and add the fresh and reconstituted-dried shitakes with 1/2 cup of the reserved mushroom water. Cook, stirring frequently, until the mushrooms are tender, about 3 minutes. Add the soy sauce and rice wine. Increase the heat to high and cook, stirring, for 1 minute. Add the remaining mushroom-soaking water.  Add the noodles to the wok and stir until heated through and coated with the sauce, about 1 minute. Garnish with the green onions and sesame seeds, if using, and serve at once.  Makes 4 servings
@@
     "Is it necessary to purchase a travel book or is it realistic that we can get similar information from other resources? Usually, most individuals have a major question on buying a travel book. So here are the pros and cons of purchasing one such book.  Advantages of a Travel Book  A travel book, which may be a paperback or e-book, comes in handy while traveling. Glancing through a travel book enables you to understand the custom and culture of a particular place in the world. So you can adapt yourself to that particular environment and stay there comfortably for longer periods.  - They Come In Handy — The travel guide comes in various forms such as, e-books, paperbacks and the file formats. You can have easy access to these books, which would assist you with all details compatible to the region you are traveling to.  - They Provide Enormous Information — Electronic or traditional travel guides provide you with answers to all types of questions such as how to learn some sayings that can be used in the place where you are traveling to? How to get data on where to reside, what to see and where to eat? How to get a clear knowledge about the history of a specific region or the atmosphere that it has?  - They Suit To Your Requirements — To access full information about a specific country or a region, both types of general and specific travel books are made available. The e-book may easily fit into your e-book reader whereas the paperback can fit into your backpack.  Disadvantages of Travel Book  - The Price — The e-book and paperback travel guides are very expensive compared to the information obtained from travel websites or from those who have moved or traveled to that region.  - Qualitative Images In Travel Books — Most travel books are in black and white. Only a few e-books consist of colored photos. Hence make a thorough revision before purchasing a travel guide or an e-book.  - Travel Books Make The Trip Less Natural — Traveling can be made more spontaneous by acquiring suggestions from locals than from travel books."
@@
In Section 4.6, we examined the results of the contamination analysis with respect to MMLU. Here, we present some analysis with respect to the other validation sets. Instead of the MMLU-specific decontamination that we performed in Section 4.6, here we follow a more general approach based on token overlaps.
@@
We examine the difference of performance of the same 7B-2x model trained on DCLM-baseline, between the full evaluation set and the evaluation samples that are marked as “not dirty” per the criterion in Touvron et al. [162] (less than 80 % of the tokens in the sample are marked as contaminated), and between the full evaluation set and samples marked as “clean” using the same criterion (less than 20 % of the tokens are marked). Results can be seen in Figure 10, where we see that the difference in performance over the full dataset and the “not dirty” samples is minimal. In fact, for BoolQ and SQuAD, which are marked as highly contaminated, our model performs slightly better on the “not dirty” subset. Moreover, in Figure 11 we see that the difference in performance between the full evaluation set and the “clean” subset is similarly small for most datasets. We note here that it’s difficult to identify a correct threshold for what counts as a contaminated sample (as 20 % token overlap might lead to many false positives, but at the same time 80 % might be too high to detect all contaminated samples).
@@
[…]
@@
We present the results in Table 26. We find that DCLM-baseline finetuned with OH-2.5 outperforms various instruct models such as Zephyr-Beta-7B and Gemma-Instruct-7B. This indicates that we can elicit high-quality responses from the pretrained DCLM-baseline model with instruction-tuning. In addition, we observe that the DCLM-baseline slightly lags behind Mistral-7B-OH-2.5 meaning DCLM-baseline is competitive with other existing models of the same scale for finetuning. The small difference in performance might be attributed to the DCLM-baseline w/ OH-2.5 having longer generations on average than Mistral-7B w/ OH-2.5 or the lesser number of tokens seen during DCLM-baseline pretraining in comparison to Mistral-7B.
A follow-up question is whether DCLM-baseline can be finetuned to be even more competitive with models of similar scale. To further improve the instruction-following capabilities of DCLM-baseline, we curate a custom dataset, DCLM-IT, by combining some of the best instruction-tuning datasets including UltraFeedback [45], Tulu-v2 SFT [76], CodeFeedback [192], OH-2.5, Nectar [195], NoRobots [132], WildChat [190], WebInstruct [184], and StarCoder2-Self-OSS-Instruct [169]. There are roughly $4$ million instances and $8$ billion tokens in this dataset. Subsequently, we perform instruction-tuning and response generation of DCLM-baseline on this dataset with the training recipe mentioned above. We present the results in Table 26. We find that DCLM-IT outperforms DCLM-baseline w/ OH-2.5 by $2.8$ percentage points. Our results highlight that there is room to enhance the instruction-following capabilities of DCLM-baseline with better datasets such as DCLM-IT. We further clarify that the current instruction-tuned models do not undergo any alignment procedures such as PPO [142], DPO [129] or others [52, 32, 16, 107]. We leave the development of aligned versions of DCLM-baseline for future research.
@@
The final run 2.5T shown in Figure 1 and Table 9 uses a two stage training procedure as followed in Groeneveld et al. [68], Hu et al. [74] and Team et al. [154]. For stage 1 we use the hyper-parameters from Table 27.
@@
After 2T tokens, we cooldown on a re-weighted pre-training distribution. For the cooldown distribution we use a mix of 70% DCLM-baseline with a tighter fastText threshold (top 7% rather than top 10%) and 30% ProofPile. We keep all the hyperparameters the same as Table 27, so we cooldown to the same final learning rate, just over a smaller number of tokens. Before the cool-down MMLU performance was approximately 52%, and the LR was approximately $1\times 10^{-3}$.
We performed 2 independent cooldowns, one for 270B tokens and another for 200B tokens, and created a “model soup" [173] with a weight of 0.8 on the 270B cooldown and a weight of 0.2 on the 200B cooldown. Thus we the total number of tokens seen by this model is 2.5T. We present results of each individual cooldown and the model soup in Table 28. The model in Figure 1 and Table 9 uses the final model soup after long-context training for 100B tokens as described in Section P.2.